# Refining Noise Mitigation in NISQ Hardware Through Qubit Error Probability

### v1

@@ 3 Using Qubit Error Probability in the Zero Noise Extrapolation, the ZEPE method
Figure 2: Transpiled quantum circuit for the Ising Hamiltonian evolution (equations 1 and 2) for a 8 qubits system. In blue are represented the basic gate operations used in $ibm\_torino$ quantum system for two (dark and light blue) trotterization steps. In red, the CZ entangle gates used for the error amplification are represented. The different color intensity accounts for each of the two error amplification factor.
@@
It is important to note that, as in ZNE, ZEPE is compatible with other error mitigation techniques. Here, we have used the T-REX mitigation technique [48, 49, 50] as implemented in Qiskit to eliminate the measurement errors. This is particularly relevant, because when we use ZEPE we cannot increase the error due to the measurement because the final state would not be invariant, meaning it cannot be mitigated within this framework. Since we have mitigated this source of error by applying T-REX, we did not take it into account while calculating the QEP, assuming it has been largely reduced.
@@
It has been shown that ZNE performs better when combined with twirling techniques [51]. Since we aim to compare our method against ZNE, we have also incorporated this error mitigation technique in our calculations. Another important issue that we have to take into account is the possible calibration problems explained above. Then, the role of warnings identifying if a qubit suffers from calibration issues or unexpected behavior is crucial, since the QEP may be miscalculated, leading to ineffective mitigation.
@@
Figure 3: Use of ZEPE to mitigate the error in calculating the magnetization of a 68 qubit system after $11$ Trotter steps in $ibm\_torino$. The green points are the raw, factor $1$, factor $2$, factor $3$ values of the magnetization as a function of the mean qubit error probability (QEP). The blue star represents the theoretical exact result, and the dashed red line is the linear error extrapolation. Finally, the green star stands for the mitigated expectation value obtained from the extrapolation to the zero-error probability.
@@
In figure 3, we calculate the magnetization of an $11$-Trotter-step Ising system as a function of the average QEP. The first green point represents the result of the circuit without ZNE or ZEPE mitigation techniques. The other three green points correspond to the results obtained after adding a set of two-qubit gates as indicated above (factor $1$, factor $2$, and factor $3$). The blue star is the theoretical exact value, and the dashed red line represents the zero error probability extrapolation. We can see that this method effectively mitigates noise mostly due to the gate errors and qubit state decay. It is important to stress that this method, like ZNE, is heuristic, so we cannot guarantee that it will be able to reduce all the noise.
@@
Figure 4: ZNE and ZEPE magnetizations as a function of the number of Trotter steps. Different ZEPE colors correspond to different extrapolation adjustments, i.e. using different set of error amplification factors. For instance, blue line labeled as (1,2,3,4,5) corresponds to the linear extrapolation performed with the raw, factor 1, factor 2, factor 3, factor 4, and factor 5 points.
@@
Finally, we have analyzed ZEPE as a function of the error amplification factor. Different factors can be used to perform the extrapolation. We have run several quantum circuits increasing the number of Trotter steps using amplification error factor ranging from 1 to 5. Then, we used different combination of the expectation values of the amplified circuits to do the error zero probability extrapolation. Figure 4 shows the results of the extrapolated values using different error amplification factor for the QEP calculation. We can see that the best mitigation corresponds to the use of ”raw” and $1$, $2$, and $3$ factors to extrapolate.
@@
Among all the extrapolation options, we have chosen to use the linear extrapolation for two reasons. First, we may expect that the error while calculating most operators, especially the magnetization, scales linearly with the mean QEP. Second, we aim to avoid extrapolations such as the exponential one, which can induce to a significant deviation from the actual expectation value.
4 The comparison between ZNE and ZEPE
@@
Once we have developed the ZEPE method, we aim to compare it with the ZNE function implemented in Qiskit. As we have previously done, we compute a 68-qubit chain in $ibm\_torino$ quantum system for different numbers of Trotter steps. ZEPE has two main advantages in comparison with ZNE. First, using ZEPE, we have enhanced control over error and noise, utilizing the quantum computer’s calibration to more precisely monitor noise behavior. Second, increasing the noise is more straightforward and affordable, as it can be achieved by adjusting the circuit depth which enable us to apply it for more complex circuits.
@@
Figure 5: Comparison on the reliability between the different error mitigation techniques. Left: Absolute value of the total magnetization as a function of the amount of Trotter steps. The raw (orange), ZNE (blue), ZEPE (green) and exact value (solid line at $M=68$) are included. Right: Relative error of the raw and extrapolated values compaired with the exact theoretical result.
@@
In figure 5, we can see the absolute magnetization (left) of the raw, ZNE, and ZEPE results, and the absolute error (right) as a function of the number of Trotter steps. We can see that our method reduces the error more effectively than ZNE. ZEPE performs better for $6$-$16$ Trotter steps. The effectiveness of the method directly depends on the entangle depth of the original quantum circuit. On the one hand, if the raw circuit has a small error probability, i.e. short depth, we cannot mitigate a significant portion of it, since the currently NISQ quantum machine is accurate enough to handle it. On the other hand, if the raw circuit has a mean QEP larger than $0.6$, the error due to circuit time and the error propagation on the gate operations starts to be sufficiently large for an effective mitigation. In this regime, ZEPE performs better than ZNE because the latter suffers of incommensurate circuit time execution which yields in a inaccurate manner to account for noise.
@@
We have shown that ZEPE can be an easily tuneable error mitigation technique. It requires minimal quantum computational resources and performs better than standard ZNE. This new method has been used on IBM quantum computers but can be extended to any type of hardware.
5 Conclusions
@@
In this work, we have investigated the primary sources of errors in quantum computers to define the QEP, a fundamental metric for quantifying the likelihood of qubit errors. Given the inherent noise and instability in NISQ devices, to accurately estimate and mitigate these errors is essential for improving quantum computation reliability. The QEP metric provides a systematic approach to assessing errors at the qubit level, offering valuable insights into error propagation and its impact on quantum circuits. By leveraging QEP, we aim to enhance the precision and effectiveness of error mitigation strategies in quantum computing.
@@
One of the most critical aspects of the NISQ era is the development of efficient error mitigation techniques to bridge the gap between noisy quantum hardware and the desired quantum advantage. In this context, we explored how QEP can be utilized to improve one of the most commonly used techniques, the zero-noise extraploation ZNE. Standard ZNE methods attempt to infer error-free results by executing quantum circuits at different noise levels and extrapolating back to a noiseless regime. However, these approaches have limitations, particularly in their sensitivity to noise models and the complexity of implementation. By integrating QEP into this process, we introduced a novel extrapolation method, ZEPE, which refines the standard ZNE approach and enhances its overall performance.
@@
The results of our study demonstrate that ZEPE offers significant advantages over traditional ZNE methods. It requires minimal quantum computational resources while achieving higher accuracy in error mitigation. The adaptability and efficiency of ZEPE makes it a promising technique for improving quantum computations in the NISQ era. Additionally, ZEPE is easily tunable, allowing users to optimize its parameters based on the specific noise characteristics of a given quantum system. These attributes make it particularly valuable for practical implementations where computational overhead must be minimized while maintaining high levels of precision.
@@
Although our research has primarily focused on IBM quantum computers, the ZEPE method is not hardware-specific and can be applied to other quantum computing platforms. This flexibility ensures that ZEPE can contribute to a wide range of quantum technologies, enhancing their error resilience and overall performance. Future work could explore further refinements to the technique, including its integration with other error mitigation methods and its application to more complex quantum algorithms. By advancing robust and efficient error mitigation strategies, such as ZEPE, we move closer to achieving practical quantum utility, where quantum computations can rival and surpass classical methods in real-world applications.

### v2

@@ III Using Qubit Error Probability in the Zero Noise Extrapolation
Figure 2: Transpiled quantum circuit for the Ising Hamiltonian evolution (equation 1) for a 8 qubits system in the Clifford point ($J=\pi$, and $h=0$). In blue are represented the basic gate operations used in $ibm\_torino$ quantum system for two (dark and light blue) trotterization steps. In red, the CZ entangle gates used for the error amplification are represented. The different color intensity accounts for each of the two error amplification factor.
Once we have defined the QEP, we can use it as the extrapolation variable in the ZNE method. The idea is the same as in standard ZNE: we execute multiple circuits with the same expectation value as the original one but ensuring that the replicas will have larger error probabilities each time. To do so, we insert pairs of controlled two-qubit native gates into all connected qubits in the circuit, which in the case of $ibm\_torino$ are controlled-Z gates (CZ). We know that a pair of these two-qubit gate is just a unitary operation, so the expectation value remains the same as the original one. The addition of this set of gates defines a factor of one. We can add as many factors as we want to further amplify the noise, and hence the error probability (see figure 2). As we pointed out before, we use the mean QEP to measure the "error", and then we extrapolate the mean QEP to zero to correct our result.
@@
It is important to note that, as in the standard ZNE, these calculations are compatible with other error mitigation techniques. Here, we have used the T-REX mitigation technique Van Den Berg, Minev, and Temme (2022); Chen et al. (2019); Bravyi et al. (2021) as implemented in Qiskit to eliminate measurement errors. This is particularly relevant because when we use ZNE we cannot increase the error due to the measurement because the final state would not be invariant, meaning it cannot be mitigated within this framework. Since we have mitigated this source of error by applying T-REX, we did not take it into account while calculating the QEP, assuming it has been largely reduced.
@@
ZNE has been shown to perform better when combined with twirling techniques Wallman and Emerson (2016). Since we aim to compare our method against standard ZNE, we have also incorporated this error mitigation technique in our calculations. Another important issue to take into account is the possible calibration problems explained above. Then, the role of warnings identifying if a qubit suffers from calibration issues or unexpected behavior is crucial, since the QEP may be miscalculated, leading to ineffective mitigation.
@@
Figure 3: Use of QEP to mitigate the error in calculating the magnetization of a 68 qubit system after $11$ Trotter steps with $\delta_{t}=1/4$, $J=\pi$, and $h=0$ in $ibm\_torino$. The green points are the raw, factor $1$, factor $2$, factor $3$ values of the magnetization as a function of the mean qubit error probability (QEP). The blue star represents the theoretical exact result, and the dashed red line is the linear error extrapolation. Finally, the green star stands for the mitigated expectation value obtained from the extrapolation to the zero-error probability.
@@
In figure 3, we calculate the magnetization of an $11$-Trotter-step Ising system as a function of the average QEP. The first green point represents the result of the circuit without any ZNE mitigation techniques. The other three green points correspond to the results obtained after adding a set of two-qubit gates as indicated above (factor $1$, factor $2$, and factor $3$). The blue star is the theoretical exact value, and the dashed red line represents the zero error probability extrapolation. We can see that this method effectively mitigates noise mostly due to the gate errors and qubit state decay. It is important to stress that this method, like standard ZNE, is heuristic, so we cannot guarantee that it will be able to reduce all the noise.
@@
Figure 4: ZNE with QEP magnetizations with $\delta_{t}=1/4$, $J=\pi$, and $h=0$ as a function of the number of Trotter steps. Different colors correspond to different extrapolation adjustments, i.e. using different set of error amplification factors. For instance, blue line labeled as (1,2,3,4,5) corresponds to the linear extrapolation performed with the raw, factor 1, factor 2, factor 3, factor 4, and factor 5 points.
@@
Finally, we have analyzed ZNE as a function of the error amplification factor. Different factors can be used to perform the extrapolation. We have run several quantum circuits increasing the number of Trotter steps using amplification error factor ranging from 1 to 5. Then, we used different combination of the expectation values of the amplified circuits to do the error zero probability extrapolation. Figure 4 shows the results of the extrapolated values using different error amplification factor for the QEP calculation. We can see that the best mitigation corresponds to the use of "raw" and $1$, $2$, and $3$ factors to extrapolate.
@@
Among all the extrapolation options, we have chosen to use the linear extrapolation for two reasons. First, we may expect that the error while calculating most operators, especially the magnetization, scales linearly with the mean QEP. Second, we aim to avoid extrapolations such as the exponential one, which can induce to a significant deviation from the actual expectation value. This is due to the fact that an small error calculating the exponent may induce a huge error in the value we are extrapolating.
IV Comparison between standard ZNE and ZNE with QEP
@@
Once we have developed the ZNE with QEP calculation method, we aim to compare it with the standard ZNE function implemented in Qiskit. To do so, we will compute the previously explained trotterization for the Ising Hamiltonian for two different cases. In the first one, we compute a 68-qubit chain in $ibm\_torino$ quantum system for different numbers of Trotter steps fixing $\delta_{t}=1/4$, $J=\pi$, and $h=0$. As we pointed before this correspond to a Clifford point so we can easily know the exact value. The second case we will fix $\delta_{t}=1/4$, $J=\pi$ and the number of Trotter steps to $15$, changing the values of $h$. This corresponds to non-Clifford points so we will do it for a $32$ qubit system in order to be able to simulate it classically. Using QEP has two main advantages in comparison with standard ZNE. First, we have enhanced control over error and noise, utilizing the quantum computer’s calibration to more precisely monitor noise behavior. Second, increasing the noise is more straightforward and affordable, as it can be achieved by adjusting the circuit depth which enable us to apply it for more complex circuits.
@@
Figure 5: Comparison on the reliability between the different error mitigation techniques. Left: Absolute value of the total magnetization as a function of the amount of Trotter steps. The raw (orange), standard ZNE (blue), ZNE with QEP (green) and exact value (solid line at $M=68$) are included. Right: Relative error of the raw and extrapolated values compaired with the exact theoretical result.
Figure 6: Comparison on the reliability between the different error mitigation techniques. Absolute value of the total magnetization as a function of the $h$ parameter. The raw (orange), standard ZNE (blue), ZNE with QEP (green) and exact value (solid line) are included.
@@
In figure 5, we can see the absolute magnetization (up) of the raw, i.e. without ZNE correction, standard ZNE, and ZNE with QEP results, and the absolute error (bottom) as a function of the number of Trotter steps. We can see that our method reduces the error more effectively than standard ZNE. It performs better for $6$-$16$ Trotter steps. The effectiveness of the method directly depends on the entangle depth of the original quantum circuit. On the one hand, if the raw circuit has a small error probability, i.e. short depth, we cannot mitigate a significant portion of it, since the currently NISQ quantum machine is accurate enough to handle it. On the other hand, if the raw circuit has a mean QEP larger than $0.6$, the error due to circuit time and the error propagation on the gate operations starts to be sufficiently large for an effective mitigation. In this regime, ZNE with QEP performs better than standard ZNE because the latter suffers of incommensurate circuit time execution which yields in an inaccurate manner to account for noise.
In figure 6, we present the absolute magnetization for the raw, standard ZNE, and ZNE with QEP results, as a function of the $h$ parameter. As in the previous case, the inclusion of QPE improves the effectiveness of standard ZNE.
@@
We have shown that the use of QPE in ZNE can be an easily tuneable error mitigation technique. It requires minimal quantum computational resources and performs better than standard ZNE. This new method has been used on IBM quantum computers but can be extended to any type of hardware.
V Conclusions
@@
In this work, we have investigated the primary sources of errors in quantum computers to define the QEP, a fundamental metric for quantifying the likelihood of qubit errors. Given the inherent noise and instability in NISQ devices, to accurately estimate and mitigate these errors is essential for improving quantum computation reliability. The QEP metric provides a systematic approach to assessing errors at the qubit level, offering valuable insights into error propagation and its impact on quantum circuits. By leveraging QEP, we aim to enhance the precision and effectiveness of quantifying the error and mitigate it in quantum computing.
@@
One of the most critical aspects of the NISQ era is the development of efficient error mitigation techniques to bridge the gap between noisy quantum hardware and the desired quantum advantage. In this context, we explored how QEP can be utilized to improve one of the most commonly used techniques, the zero-noise extraploation ZNE. Standard ZNE methods attempt to infer error-free results by executing quantum circuits at different noise levels and extrapolating back to a noiseless regime. However, these approaches have limitations, particularly in their sensitivity to noise models and the complexity of implementation. By integrating QEP into this process, we improve the standard ZNE approach and enhance its overall performance.
@@
The results of our study demonstrate that the inclusion of QPE offers significant advantages over traditional ZNE methods. As standard ZNE, and in contrast with other mitigation techniques, it requires minimal quantum computational resources while achieving higher accuracy in error mitigation. Additionally, it is easily tunable, allowing users to optimize its parameters based on the specific noise characteristics of a given quantum system. These attributes make it particularly valuable for practical implementations where computational overhead must be minimized while maintaining high levels of precision.
@@
Although our research has primarily focused on IBM quantum computers, the method is not hardware-specific and can be applied to other quantum computing platforms. This flexibility ensures that it can contribute to a wide range of quantum technologies, controlling their errors and enhancing their error resilience and overall performance through the improvement of ZNE. Future work could explore further refinements to the technique, including its integration with other error mitigation methods and its application to more complex quantum algorithms. By advancing robust and efficient error mitigation strategies, we move closer to achieving practical quantum utility, where quantum computations can rival and surpass classical methods in real-world applications.
@@ Acknowledgments
Author Declarations
Conflict of Interest
The authors have no conflicts to disclose.
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.